OpenStudy (anonymous):

when you are solving for a variable in an inequality do you flip the inequality sign if you divide?

OpenStudy (anonymous):

i like flipping in all circumstances but no only if the number is negativef

OpenStudy (anonymous):

so when I divide by P in this equation I dont flip the sign?

OpenStudy (calculusfunctions):

You reverse the inequality sign whenever i). you multiply or divide both sides of the inequation by a negative, ii). read the inequation backwards, or iii). take the reciprocal of both sides as long as both sides are positive or both sides are negative.

OpenStudy (anonymous):

god i want sooo bad to know where this question comes from please please please i beg you to tell me

OpenStudy (anonymous):

flvs why?

OpenStudy (anonymous):

how did i guess?

OpenStudy (anonymous):

you cannot solve this problem for n it cannot be done and you asked a very good question, when do you flip the inequality

OpenStudy (anonymous):

the answer is when you multiply or divide by a negative number but the idiot who wrote this question does not understand that a variable can either be positive or negative

OpenStudy (anonymous):

The Algebra teachers at my school are terrible so I take it online. I've come across a number of questions like this, put I've found even more in the school text book.

OpenStudy (anonymous):

please can you give me the id number of this question ? it is shameful

OpenStudy (anonymous):

I am not solving for n. I am finding how I would solve for n if there where numbers in place of the variables. Would you like to see the whole question?

OpenStudy (anonymous):

i once wrote to FLVS asking why their questions sucked so bad (not in so many words) they asked for id numbers if i had it i could give an example

OpenStudy (anonymous):

i would love to

OpenStudy (anonymous):

what do you mean by "id numbers"?

OpenStudy (anonymous):

the questions come with id numbers, maybe you do not see them the image will be good enough they are criminals

OpenStudy (anonymous):

about the original question, I wouldn't flip the sign?

OpenStudy (anonymous):

here is the thing YOU DON'T KNOW!! no one knows, because you don't know if the variables are positive or not you can't know but we can still pretend

OpenStudy (anonymous):

the variable is negative

OpenStudy (anonymous):

no it isn't

OpenStudy (anonymous):

it is a variable

OpenStudy (anonymous):

it is look at the picture. it says -np-4

OpenStudy (anonymous):

god help me it is not your fault

OpenStudy (jhannybean):

If you were not given answer choices, would you be able to determine whether p or n were negative or positive?

OpenStudy (anonymous):

lets go ahead and solve

OpenStudy (anonymous):

since it says -np I should think so

OpenStudy (anonymous):

NNOOOOO that is the whole point!

OpenStudy (anonymous):

ok. I thought the answer would be either a or c

OpenStudy (anonymous):

you are being lead to believe that \(-p\) is negative because it has a minus sign

OpenStudy (anonymous):

i understand this, but you should be smarter than the florid school board if for example \(p=-4\) then \(-p=4\)

OpenStudy (anonymous):

lets solve then i will go take a valium

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

\[-np-4\leq 2(c-3)\\ -np-4\leq 2c-6\\\ -np\leq 2c-2\]

OpenStudy (anonymous):

ok then you divide by p

OpenStudy (anonymous):

now they want you do divide by \(-p\) and change the inequality because they are so uneducated as to think \(-p\) is negative and get \[n\geq -\frac{2c-2}{p}\]

OpenStudy (anonymous):

ok so th3 answer is c?

OpenStudy (anonymous):

sorry, the, old habits

OpenStudy (anonymous):

that is what they want you to think, yes

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

write to them and tell them that they don't understand what a variable is

OpenStudy (anonymous):

that they have no business teaching math

OpenStudy (anonymous):

i know i will

OpenStudy (anonymous):

it's how they've taught it to me...

OpenStudy (anonymous):

thanks for the screenshot btw

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

they should be prosecuted

OpenStudy (anonymous):

feel free. XD